Points A, B, and C are collinear. Point B is between A and C. Find the length indicated.
1) BC=2x+23, AC=x+25, and AB=10. Find BC

Answer by stanbon(75887)

Points A, B, and C are collinear. Point B is between A and C. Find the length indicated.
1) BC=2x+23, AC=x+25, and AB=10. Find BC
--------------
A......10........B.......2x+23.........C
=================
AC = AB + BC
------
x+25 = 10 + 2x+23
-------
x = -8
-------
BC == 2x+23 = 2*-8+23 = -16+23 = 7
=====================================
Cheers,
Stan H.
===========